Data suggest that 232 male deaths from lightning strikes and 55 female deaths from lightning strikes occur... Assume that for a randomly selected lightning death there is a 0.8 probability that the victim is a male

A) find the probability that three random people killed by lightning strikes are all male

B) find the probability that three random people killed by lightning strikes are all female

C) Find the probability that among three people killed by lightning stirkes atleast one is a male

D) If five people killed by lightning strikes are randomly selected, find the probability that exactly three of them are males.

E) A study involves random selection of different groups of 50 people killed by lightning strikes. for those groups find the mean and standard deviation for the number of male victims.
Answer #1

A)

0.8^3 = 0.512

B)

0.2^3 = 0.008

C)

P(atleast 1 male) = 1- P(all female) = 1-0.008 = 0.992

D)

P(x=3) = C(5,3)*(0.8^3)*(0.2^2) = 0.2048

E)

Mean = n*p = 50*0.8 = 40

Standard dev = (npq)^0.5 = (50*0.8*0.2)^0.5 =2.828